OpenStudy (anonymous):

find slope of the line that goes through the given points ... (-2,1) and (1/4,5)

OpenStudy (anonymous):

slope = (-2-5)/(1-(1/4) = -7 /(3/4) =-28/3

OpenStudy (anonymous):

that's not an option as an answer..

OpenStudy (angela210793):

 where x1;y1 r the coordinates of the 1st point and x2;y2 the coordinates of the 2nd point